An experiment consists of choosing one card from a standard deck and then replacing it. The experiment was done several times and the results are :8 hearts, 8 diamonds, 6 spades, and 6 clubs. Find the experimental probability that a card is red

Answer by jim_thompson5910(35256)

In a standard deck, the red cards are hearts and diamonds. There are "8 hearts, 8 diamonds", so 8+8 = 16 red cards.

There are "8 hearts, 8 diamonds, 6 spades, and 6 clubs", so 8+8+6+6 = 28 cards total

The experimental probability is the number of things we want over the number of things total, so...


experimental probability = (# of things we want)/(# of things total)

experimental probability that card is red = (# of red cards)/(# of cards total)

experimental probability that card is red = 16/28

experimental probability that card is red = 4/7 ... make sure to reduce all fractions


So the answer is 4/7